OpenStudy (anonymous):

y=x^2-2x-3 x=intercepts?

Parth (parthkohli):

Set y = 0 to get the x intercept. \( \color{Black}{\Rightarrow 0 = x^2 - 2x - 3 }\)

Parth (parthkohli):

This is more like a quadratic equation. Can you solve?

OpenStudy (mimi_x3):

Trimonials! \[x^2+x(a+b)+ab=(x+a)(x+b) \]

OpenStudy (anonymous):

x-intec 1?

OpenStudy (anonymous):

0,1?

OpenStudy (anonymous):

-1,3

OpenStudy (anonymous):

y=(x-3)*(x+1)===> I thk x=-1==>3